A car is to travel two miles. It goes the first mile at an average velocity of 30 mph. the driver wishes to average 60 mph for the entire two-mile trip. is this possible? explain.
Answer by macston(5194)

.
To average 60 mph over 2 miles you must make the entire trip in:
2miles/(60mi/hr)=1/30 hr =2 minutes
To travel 1 mile at 30 mph takes:
1mi/30mph=1/30 hr=2 minutes
You cannot make the entire 2 mile trip in two minutes because it took 2 minutes to travel the first mile, so you cannot average 60 miles per hour.